OpenStudy (anonymous):

Find y′ by implicit differentiation. e^(x/y)=3x-4y

OpenStudy (anonymous):

find the y'

OpenStudy (kainui):

Show us your best guess first. Then we'll help.

OpenStudy (kainui):

Right @mathmale ?

OpenStudy (mathmale):

Suggestion: Apply the derivative operator, (d/dx), to both sides of your equation. Treat x as the independent variable and y as the dependent variable (that is, act as though y were a function of x, which requires application of the chain rule. It'd help you, as well as possible helpers, if you'd please share your first tentative efforts towards finding the derivative.

OpenStudy (anonymous):

i'm guessing the only thing troubling you is the e^(x/y) ? :P

OpenStudy (anonymous):

uh, I dont get how to derivative the e^x/y

OpenStudy (anonymous):

yep, you are right.@iambatman

OpenStudy (mathmale):

Hints: (d/dx)e^x= e^x (d/dx)e^u = e^u * (du/dx) (Chain rule)

OpenStudy (anonymous):

^^

OpenStudy (mathmale):

That exponent, (x/y) is itself a function, and in particular is a quotient. Thus when you encounter x/y as the exponent, you'll need to apply the chain rule, which in turn requires that you apply the quotient rule. Mind giving this a stab?

OpenStudy (anonymous):

sure, give me a sec,

OpenStudy (anonymous):

I don't know if this will help you but, what are the derivatives of e^2x and e^(2/x)?

OpenStudy (anonymous):

err e^(x/2)*

OpenStudy (anonymous):

@iambatman e^2x=2e^x, e^(x/2)=same, is it right?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

It's just chain rule, so what's the derivative of e^2x?

OpenStudy (anonymous):

What would you say? e^x = e^x, so now if you have e^2x and apply the chain rule, you get?

OpenStudy (anonymous):

so it would be e^2x=2e^2x

OpenStudy (anonymous):

Exactly, so what's the derivative of e^(x/2)?

OpenStudy (anonymous):

is it 1/2e^(x/2)

OpenStudy (anonymous):

:), so what's the derivative of e^(x/y)?

OpenStudy (anonymous):

is this right ? x/ye^x/y

OpenStudy (anonymous):

ohh, so it would be 1/ye^x/y ?

OpenStudy (anonymous):

What did you get as your final answer?

OpenStudy (anonymous):

e^x/y=1-y^2 is this right?

OpenStudy (anonymous):

d(x)3x-4y=1-y right?

OpenStudy (anonymous):

i am still confused.....

OpenStudy (anonymous):

1-y?

OpenStudy (anonymous):

How'd you get that

OpenStudy (anonymous):

bc 3 is constant and the x is 1

OpenStudy (anonymous):

What's the derivative of 3x?

OpenStudy (anonymous):

if 3 is constant, that means it's 0

OpenStudy (anonymous):

Ok, but what's the derivative of 3x?

OpenStudy (anonymous):

Oh I see what you're saying but, no that's not how it works. I think you need to go over some derivative rules. http://www.mathsisfun.com/calculus/derivatives-rules.html

OpenStudy (anonymous):

Derivative of 3x = 3

OpenStudy (mathmale):

The exponent of " e " in the original equation is itself a function:\[\frac{ x }{ y }.\] I strongly suggest you focus on this x/y for the time being. Use the Quotient Rule and Chain Rule to find the derivative (with respect to x) of this x/y. Once you have this done, use the result in writing the derivative with respect to x of \[e ^{\frac{ x }{ y }}\]

OpenStudy (anonymous):

Sorry I didn't make this clear, I was just showing you how to get the basic derivatives, and once you figured that out, we could move to the actual problem itself :P. As @mathmale explained, that's how you would do it.

OpenStudy (anonymous):

So, what did you get as your derivative for e^(x/y)?

OpenStudy (anonymous):

1/y(e^x/y)

OpenStudy (anonymous):

Use the chain rule

OpenStudy (anonymous):

e^(x/y)*(x/y)'

OpenStudy (anonymous):

ohhh, then i should use quotient on (x/y)' right?

OpenStudy (anonymous):

Yes :P

OpenStudy (anonymous):

(e^x/y)*(y-y')/y^2 is this right?

OpenStudy (anonymous):